Can anyone please help me make sense of this?? Thank you so much for your time!
Assuming that a, b and a/b(a divided by b) are whole numbers and b is not equal to 0. Explain why (a / b) xb= a ((a divided by b) x b = a).

Answer by stanbon(75887)

Assuming that a, b and a/b(a divided by b) are whole numbers and b is not equal to 0. Explain why (a / b) xb= a ((a divided by b) x b = a).
------------------
If a|b then b = ka where k is a whole number
and k = b/a
----------------
So, (a/b)x b
= (a/ka)x b
= (1/k)x b
= (1/(b/a))x b
= (a/b)x b
= a
===================================================================
Cheers,
Stan H.